OpenStudy (anonymous):

16x(x-2)=8x-25 solve by factoring

OpenStudy (anonymous):

16x(x-2)=8x-25 16x^2 - 32x=8x-25 16x^2-40x+25=0

OpenStudy (anonymous):

Now, can u factor it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

16x^2-40x+25=0 (4x)^2-2*4x*5+(5)^2=0

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Welcom